OpenStudy (anonymous):

One side of a triangle is 15 inches, and the area of the triangle is 90 sq. inches. Find the area of a similar triangle in which the corresponding side is 9 inches. Can someone help me please i dnt have time cus today is my last day of school so someone please help ill b ur fan ill give u medals

OpenStudy (id21):

32,4 in^2

OpenStudy (anonymous):

1. The areas of the similar triangles will be in the proportion of squares of the ratio of any 2 corresponding sides. Let the required area be x squared inches. Then 90 / x = ( 15 / 9 )^2 ( Here (15/9)^2 means square of 15/9 ) Therefore x = 90 / ( 5/3)^2 = 90 * 9 / (5 * 5) = 18 * 9 / 5 = 162 / 5 = 32.4 sq. inches. ( '*' denotes multiplication ) [ this may help u ]

OpenStudy (id21):

I think that she don't care about theory, just need an answer.

OpenStudy (anonymous):

hey thats not true

OpenStudy (id21):

But most of the problems you write is so similar. You must try to solve it by yourself. Because it is not hard, it is easy if you'd read the answers with explanations, which everyone writes to you.

OpenStudy (anonymous):

well i try to but i have a learning disablilty which is hard for me to understand im not like those ppl who can get it right then